OpenStudy (anonymous):

A system of equations is shown below: x + 3y = 5 (equation 1) 7x − 8y = 6 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 3x + y = 5 and 8x −7y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 8x − 5y = 11 and 7x − 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 15x + 13y = 17 and 7x − 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations −13x + 15y = 17 and 7x − 8y = 6 is the same as the solution to the given system of equations

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

will award medals

OpenStudy (anonymous):

brb gtg eat

OpenStudy (welshfella):

add the 2 equations term by term and replace the first equation by the result

OpenStudy (anonymous):

8x-5y=11

OpenStudy (anonymous):

is B the answer